Lead, Pb, is used in solders and as radiation shield. It adopts the face centered cubic unit structure with each of the atom having a diameter of 0.35 nanometer. Find the theoretical density of lead if the molecular weight of this metal is 207.2 g/mol.


1
Expert's answer
2020-10-01T07:51:30-0400

For FCC, N=4 , r= 0.35/2 x 10^(-6) mm , A=207.2 g/mole given..

edge length ,we know , a=4r/Sqrt(2) = 4.95x10^(-8) cm

hence density (roh) = (NA)/{ 6.023x10^(23) * a^3 } = 11.35 g/cm^3.

theoretical density of lead = 11.35 g/cm^3